# Fast Policy Synthesis with Variable Noise Diffusion Models

### v3

@@ III-A Action Buffer
Similar to Diffusion Policy [1] we denoise an clean action sequence $\mathbf{A}_{t}$ of $T_{a}$ future actions. However, in contrast to Diffusion Policy, our proposed method SDP denoises each future action in $\mathbf{A}_{t}$ with a different level of increasing noise. Actions outside the immediate action that is executed is stored in action buffer that is reused to generate actions in future observations.
@@
In SDP, the action $\mathbf{A}_{t,i}$ at action timestep $i$ is denoised at a separate noise level $k_{i}$, where:
$\mathbf{A}^{k_{i}-1}_{t,i}\leftarrow\alpha_{k_{i}}(\mathbf{A}^{k_{i}}_{t,i}-\gamma_{k_{i}}\epsilon_{\theta}(\mathbf{A}_{t,i}^{k_{i}};\mathbf{O}_{t},\mathbf{k})_{i}+\mathcal{N}(0,\sigma_{k_{i}}^{2})).$ (2)
Here, $\epsilon_{\theta}(\mathbf{A}_{t,i}^{k_{i}};\mathbf{O}_{t},\mathbf{k})_{i}$ denotes the element at index $i$ in the noise prediction. Each action starts at a diffusion level of $k_{i}=N$ corresponding to pure noise with $k_{i}=0$ corresponding to a clean action. We store future actions $\mathbf{A}_{t,i}$ for $i>1$ in an action buffer that is used to generate actions in future observations (illustrated in Figure 2).
@@ IV Experiments
Our experiments answer three questions: (a) What choice of action primer and noise injection schemes leads to the highest performance?, (b) What is the reduction in sampling speed for SDP relative to standard diffusion-based policies? and (c) How does Streaming Diffusion Policy compare to baseline methods on simulated and real robotic tasks?

### v4

@@ III-A Action Buffer
Similar to Diffusion Policy [1] we denoise an clean action sequence $\mathbf{A}_{t}$ of $T_{a}$ future actions. However, in contrast to Diffusion Policy, our proposed method SDP denoises each future action in $\mathbf{A}_{t}$ with a different level of increasing noise. Actions outside the immediate action that is executed are stored in action buffer that is reused to generate actions in future observations.
@@
In SDP, the action $\mathbf{A}_{t,i}$ at action timestep $i$ is denoised at a separate noise level $k_{i}$ from $\mathbf{k}=[k_{0},k_{1},\dots,k_{T_{a}-1}]$:
$\mathbf{A}^{\mathbf{k}-1}_{t}\leftarrow\alpha_{\mathbf{k}}(\mathbf{A}^{\mathbf{k}}_{t}-\gamma_{\mathbf{k}}\epsilon_{\theta}(\mathbf{A}_{t}^{\mathbf{k}};\mathbf{O}_{t},\mathbf{k})+\mathcal{N}(0,\sigma_{\mathbf{k}}^{2})),$ (2)
with $\alpha_{\mathbf{k}}=[\alpha_{k_{0}},\alpha_{k_{1}},\dots,\alpha_{k_{T_{a}-1}}]$. The coefficients $\gamma_{\mathbf{k}},\sigma_{\mathbf{k}}$ are constructed similarly. Each action starts at a diffusion level of $k_{i}=N$ corresponding to pure noise with $k_{i}=0$ corresponding to a clean action. We store future actions $\mathbf{A}_{t,i}$ for $i>1$ in an action buffer that is used to generate actions in future observations (illustrated in Figure 2).
@@ IV Experiments
Our experiments answer three questions: (a) What choice of action primer and noise injection schemes leads to the highest performance?, (b) What is the increase in sampling speed for SDP relative to standard diffusion-based policies? and (c) How does Streaming Diffusion Policy compare to baseline methods on simulated and real robotic tasks?